# Information Propagation in Predator-Prey Dynamics of Turbulent Plasma

### v1

@@ Abstract
Magnetically confined fusion plasmas exhibit predator-prey-like cyclic oscillations through the self-regulating interaction between drift-wave turbulence and zonal flow. To elucidate the detailed mechanism and causality underlying this phenomenon, we construct a simple stochastic predator-prey model that incorporates intrinsic fluctuations and analyze its statistical properties from an information-theoretic perspective. We first show that the model exhibits persistent fluctuating cyclic oscillations called quasi-cycles due to amplification of intrinsic noise. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles but quasi-cycles, and that such quasi-cycles may be widely observed under various conditions. For this model, we further prove that information of zonal flow is propagated to turbulence. This information-theoretic analysis may provide a theoretical basis for regulating turbulence by controlling zonal flow.
@@ Introduction.
—Predator-prey dynamics can be observed ubiquitously in various systems such as ecosystems [1, 2, 3], genetic systems [4], fluids [5, 6, 7, 8], or even quantum systems [9]. Magnetically confined fusion plasmas also exhibit predator-prey-like cyclic oscillations in various situations, including the self-regulating process between zonal flow and drift-wave turbulence [10, 11, 12, 13], the intermediate phase (I-phase) during the low-to-high confinement (L-H) transition [14, 15, 16, 17, 18], the formation and collapse of internal transport barriers [19], and self-regulated oscillations of magnetic islands [20]. Among these, the interplay between zonal flow (predator) and drift-wave turbulence (prey) is believed to play a crucial role in suppressing anomalous heat and particle transport, and therefore, intensive attempts have been made to regulate turbulence by controlling zonal flow. Despite this importance, the detailed mechanism and causality underlying the cyclic oscillations have not been elucidated [12]. To deepen our understanding of the self-regulating process between zonal flow and turbulence, it is desirable to construct a minimal model that exhibits the cyclic oscillations and investigate the causality within the oscillations.
@@
The main aim of this Letter is twofold. The first is to construct a simple model for the cyclic oscillations observed in plasma turbulence. Note that the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which incorporates the self-regulation mechanism between turbulence and zonal flow, does not exhibit a stable cyclic oscillation. To address this issue, we aim to construct a stochastic predator-prey model that incorporates the effects of intrinsic noise from the perspective of statistical physics. Since it has been pointed out in the field of ecology that it is essential to consider intrinsic fluctuations in predator-prey dynamics [22, 23], we expect that such a statistical physics approach provides useful insights into plasma turbulence. The second is to quantify the statistical causality between zonal flow and turbulence in this model through the lens of information theory. While several previous studies have attempted to quantify the statistical causality in plasma turbulence using information-theoretic quantities, such as (net) transfer entropy [24, 25] or information length [26, 27, 28, 29, 30, 31, 32], these quantities do not change sign under time reversal and thus can be nonzero even in an equilibrium state where all probability currents vanish. Instead, here we employ information flow, which is a pivotal concept within the framework of information thermodynamics [33, 34, 35, 36, 35, 37, 38] and has recently been applied to standard fluid turbulence [39, 40, 41].
@@
The constructed model can be interpreted as a standard two-variable predator-prey model that takes into account intrinsic fluctuations. We show that the model exhibits persistent fluctuating cyclic oscillations, called quasi-cycles [22, 23], due to stochastic amplification. The quasi-cycles can be observed for a wide range of parameters, including parameter values consistent with experiments [14] and gyrokinetic simulations [42]. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles, but rather quasi-cycles. Our analysis of the resonance condition also implies that such quasi-cycles may be widely observed under various conditions.
@@
For this model, we quantify the causality between zonal flow and turbulence using information flow. Specifically, we analytically show that the information on zonal flow is propagated to turbulence. This information propagation occurs as long as zonal flow and turbulence coexist, regardless of the presence or absence of quasi-cycles. Thus, our information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow.
@@
[…]
System size expansion.
—Since we are interested in the collective behavior of the system, we consider coarse-grained descriptions by taking the infinite volume limit $\Omega\rightarrow\infty$. Here, we note that the two limits $t\rightarrow\infty$ and $\Omega\rightarrow\infty$ do not commute in this model, which is known as Keizer’s paradox [44, 45]. Indeed, if we take the limit $t\rightarrow\infty$ first, then the probability distribution converges to a unique stationary distribution $p_{\mathrm{ss}}({\bm{n}})=\delta_{{\bm{n}},{\bm{0}}}$, which implies that eventually there will be no zonal flows and turbulences. Below, we focus on the opposite case, where we first take $\Omega\rightarrow\infty$ and then $t\rightarrow\infty$. The coarse-grained dynamics in the limit $\Omega\rightarrow\infty$ can be systematically obtained by applying van Kampen’s system size expansion [46]. To separate the fluctuating part from the average part, we introduce a new stochastic variable ${\bm{r}}_{t}=(z_{t},w_{t})\in\mathbb{R}^{2}$ as ${\bm{n}}_{t}/\Omega={\bm{c}}_{t}+{\bm{r}}_{t}/\Omega^{1/2}$, where ${\bm{c}}_{t}=(c^{Z}_{t},c^{W}_{t})\in\mathbb{R}^{2}_{\geq 0}$ denotes the mean energy density to be determined. By substituting this expression into Eq. (1) and expanding in terms of $\Omega^{-1/2}$, we find that the leading order $O(\Omega^{1/2})$ yields the time evolution equation for the average part ${\bm{c}}_{t}$, which can be regarded as a coarse-grained dynamics of Eq. (1):
@@
See Ref. [47] for the detailed derivation. This equation corresponds to the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which has the same form as the Lotka–Volterra equation. There are two regimes in this model (Fig. 1(b)). When $bp<d\gamma$, the stable fixed point is $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=(0,b/\gamma)$, which represents a state without zonal flows. In contrast, when $bp>d\gamma$, the stable fixed point is $((bp-d\gamma)/p^{2},d/p)$, which represents a coexistence state. Importantly, this equation does not exhibit limit cycles, contrary to the naive expectation of the existence of predator-prey oscillations. While cyclic oscillations emerge when $\gamma=0$, these cycles are not limit cycles but structurally unstable.
Note that the coarse-grained description [Eqs. (3) and (4)] ignores intrinsic noise that is inevitably present in the original individual-level description (1). For a realistic situation where $\Omega$ is large but finite, it becomes essential to consider the intrinsic noise. To investigate the effect of the intrinsic noise on the coexistence state, we focus on the regime $bp>d\gamma$ and proceed to the subleading order $O(\Omega^{0})$ of the system size expansion by substituting $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=((bp-d\gamma)/p^{2},d/p)$. The resulting equation describes the time evolution of the fluctuating part ${\bm{r}}_{t}$ around ${\bm{c}}_{t}={\bm{c}}_{\mathrm{ss}}$ [47]:
@@
where $\tau:=bt$ denotes the dimensionless time, $\beta:=d/b$ denotes the dimensionless damping rate, and $\epsilon:=1-d\gamma/bp\in(0,1)$ denotes the ratio of time scales for $z$ and $w$. The terms $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ denote the zero-mean white Gaussian noise that satisfies $\langle\zeta^{Z}_{\tau}\zeta^{Z}_{\tau^{\prime}}\rangle=2D^{Z}\delta(\tau-\tau^{\prime})$, $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$, and $\langle\zeta^{W}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=2D^{W}\delta(\tau-\tau^{\prime})$, where $D^{Z}:=\epsilon c^{W}_{\mathrm{ss}}=\epsilon d/p$ and $D^{W}:=c^{W}_{\mathrm{ss}}=d/p$. Note that this linear Langevin equation has a form in which noise is added to the linearized Lotka–Volterra equation around the coexistence fixed point. We emphasize that the noise intensities $D^{Z}$ and $D^{W}$ are uniquely determined from the master equation (1). In other words, the noise driving the system is not artificially added but rather arises from intrinsic stochasticity. This property marks a stark difference from the stochastic models used in Refs. [26, 27, 28, 29, 30, 31, 32], where noise is artificially added to deterministic predator-prey models. The Langevin equations (5) and (6) are equivalent to the Fokker–Planck equation $\partial_{\tau}p_{\tau}({\bm{r}})=-\partial_{z}J^{Z}_{\tau}({\bm{r}})-\partial_{w}J^{W}_{\tau}({\bm{r}})$, where $p_{\tau}({\bm{r}})$ denotes the probability density of finding ${\bm{r}}_{\tau}={\bm{r}}$ and $J^{Z}_{\tau}({\bm{r}})$ and $J^{W}_{\tau}({\bm{r}})$ denote the probability currents [47].
Quasi-cycles induced by stochastic amplification.
@@
—Although the coarse-grained deterministic description [Eqs.(3) and (4)] cannot predict stable cycles, the stochastic description [Eqs. (5) and (6)] predicts persistent cyclic oscillations called quasi-cycles. Figure 1(c) shows an example of such persistent fluctuating oscillations observed in the linear Langevin equation. Notably, the fluctuations of zonal flow follow those of turbulence with a phase lag $\sim\pi/2$, which are similar to the predator-prey-like oscillations observed in a toroidal plasma [14, 15, 16, 17, 12]. This oscillatory behavior occurs when the coexistence state is a stable spiral fixed point in the coarse-grained deterministic description. More precisely, from the calculation of the power spectral density [47], we can show that the resonant stochastic amplification occurs for a wide range of parameters, where the characteristic resonant frequency $\omega_{c}:=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}$ takes a real value (see also Fig. 2(b)). Because the white noise in the Langevin equation covers all frequencies, the corresponding cycle with the resonant frequency $\omega_{c}$ is amplified. This mechanism is called stochastic amplification to distinguish it from stochastic resonance [22].
@@
The quasi-cycles differ from limit cycles in that they are stochastic oscillations where the different frequencies near $\omega_{c}$ are all excited [23, 48]. Furthermore, because quasi-cycles can occur as long as the system has a stable spiral fixed point and noise, they may be widely observed in plasma turbulence under various conditions. Although we have emphasized that our model is a qualitative description, here we present a set of parameter values consistent with experiments and simulations for reference; By comparing with the gyrokinetic simulations [42], they can be estimated as $\epsilon\sim 1$ and $\beta\sim 10^{-4}$, which satisfies the resonance condition, and the corresponding cycle frequency reads $\omega_{c}\sim 10^{-2}$ in dimensionless units.
Information propagation between zonal flow and turbulence.
—The presence of quasi-cycles suggests that intrinsic fluctuations are essential for understanding the dynamics of zonal flow and turbulence. Therefore, we scrutinize the statistical causality underlying the intrinsic fluctuations from an information-theoretic viewpoint.

### v2

@@ Abstract
Magnetically confined fusion plasmas exhibit predator-prey-like cyclic oscillations through the self-regulating interaction between drift-wave turbulence and zonal flow. To elucidate the detailed mechanism and causality underlying this phenomenon, we construct a simple stochastic predator-prey model that incorporates intrinsic fluctuations and analyze its statistical properties from an information-theoretic perspective. We first show that the model exhibits persistent fluctuating cyclic oscillations called quasi-cycles due to amplification of intrinsic noise. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles but quasi-cycles, and that such quasi-cycles may be widely observed under various conditions. For this model, we further prove that information of the zonal flow is propagated to turbulence. This result suggests that turbulence behavior may be predictable to a certain extent based on zonal flow characteristics.
@@ I Introduction
Predator-prey dynamics can be observed ubiquitously in various systems such as ecosystems [1, 2, 3], genetic systems [4], fluids [5, 6, 7, 8], or even quantum systems [9]. Magnetically confined fusion plasmas also exhibit predator-prey-like cyclic oscillations in various situations, including the self-regulating process between zonal flow and drift-wave turbulence [10, 11, 12, 13], the intermediate phase (I-phase) during the low-to-high confinement (L-H) transition [14, 15, 16, 17, 18], the formation and collapse of internal transport barriers [19], and self-regulated oscillations of magnetic islands [20]. Among these, the interplay between zonal flow (predator) and drift-wave turbulence (prey) is believed to play a crucial role in suppressing anomalous heat and particle transport, and therefore, intensive attempts have been made to regulate turbulence by controlling zonal flow. Despite this importance, the detailed mechanism and causality underlying the cyclic oscillations have not been elucidated [12]. To deepen our understanding of the self-regulating process between zonal flow and turbulence, it is desirable to construct a minimal model that exhibits the cyclic oscillations and investigate the causality within the oscillations.
@@
The main aim of this Letter is twofold. The first is to construct a simple qualitative model for the cyclic oscillations observed in plasma turbulence. Note that the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which incorporates the self-regulation mechanism between turbulence and zonal flow, does not exhibit a stable cyclic oscillation. To address this issue, we aim to construct a stochastic predator-prey model that incorporates the effects of intrinsic noise from the perspective of statistical physics. Since it has been pointed out in the field of ecology that it is essential to consider intrinsic fluctuations in predator-prey dynamics [22, 23], we expect that such a statistical physics approach provides valuable insights into plasma turbulence. The second is to quantify the predictive causality between zonal flow and turbulence in this model through the lens of information theory. Several previous studies have attempted to quantify the predictive causality in plasma turbulence using information-theoretic quantities, such as (net) transfer entropy [24, 25] or information length [26, 27, 28, 29, 30, 31, 32, 33]. However, because these quantities are not antisymmetric under time reversal, it is unclear whether they appropriately quantify predictive causality. Instead, here we employ information flow, which is a pivotal concept within the framework of information thermodynamics [34, 35, 36, 37, 36, 38, 39] and has recently been applied to various systems, including standard fluid turbulence [40, 41, 42]. Because the information flow is antisymmetric under time reversal and becomes zero in an equilibrium state where all probability currents are absent, it may reflect some aspects of the predictive causality consistent with the behavior of the probability currents.
@@
The constructed model can be interpreted as a standard two-variable predator-prey model that takes into account intrinsic fluctuations. We show that the model exhibits persistent fluctuating cyclic oscillations, called quasi-cycles [22, 23], due to stochastic amplification. Quasi-cycles are distinguished from both deterministic and stochastic limit cycles by their shorter correlation time scales [43, 44]. Furthermore, they can readily emerge in any system possessing a stable spiral fixed point subject to noise. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles, but rather quasi-cycles. Our analysis of the resonance condition also implies that such quasi-cycles may be widely observed under various conditions.
@@
For this model, we quantify the causality between zonal flow and turbulence using information flow. Specifically, we analytically show that the information on zonal flow is propagated to turbulence. This information propagation occurs as long as zonal flow and turbulence coexist, regardless of the presence or absence of quasi-cycles. Our finding suggests that turbulence behavior may be predictable to a certain extent based on zonal flow characteristics.
@@ II Setup
where the stoichiometric coefficients ${\bm{\nu}}_{\rho}=(\nu^{Z}_{\rho},\nu^{W}_{\rho})$ denote the number of reactants Z and W involved in each reaction $\rho$, and $\Omega$ denotes the system size, which may be interpreted as a total volume of the plasma. Although we can estimate parameter values that are consistent with experiments and simulations, we consider this model as a qualitative description and will not explore its quantitative aspects in depth. We note that similar models have been used for the laminar-turbulence transition in standard fluids to investigate the connection with the directed percolation universality class [5, 6, 7, 8].
III System size expansion
Since we are interested in the collective behavior of the system, we consider coarse-grained descriptions by taking the infinite volume limit $\Omega\rightarrow\infty$. Here, we note that the two limits $t\rightarrow\infty$ and $\Omega\rightarrow\infty$ do not commute in this model, which is known as Keizer’s paradox [46, 47]. Indeed, if we take the limit $t\rightarrow\infty$ first, then the probability distribution converges to a unique stationary distribution $p_{\mathrm{ss}}({\bm{n}})=\delta_{{\bm{n}},{\bm{0}}}$, which implies that eventually there will be no zonal flows and turbulences. Below, we focus on the opposite case, where we first take $\Omega\rightarrow\infty$ and then $t\rightarrow\infty$. The coarse-grained dynamics in the limit $\Omega\rightarrow\infty$ can be systematically obtained by applying the van Kampen system size expansion (also known as the linear noise approximation) [48]. From the central limit theorem, we first note that $p_{t}({\bm{n}}/\Omega)$ can be reasonably approximated as a Gaussian with a finite width of order $\Omega^{-1/2}$. To separate the fluctuating part from the average part, we introduce a new stochastic variable ${\bm{r}}_{t}=(z_{t},w_{t})\in\mathbb{R}^{2}$ as ${\bm{n}}_{t}/\Omega={\bm{c}}_{t}+{\bm{r}}_{t}/\Omega^{1/2}$, where ${\bm{c}}_{t}=(c^{Z}_{t},c^{W}_{t})\in\mathbb{R}^{2}_{\geq 0}$ denotes the mean energy density to be determined. By substituting this expression into Eq. (1) and expanding in terms of $\Omega^{-1/2}$, we can obtain the time evolution equations for ${\bm{c}}_{t}$ and ${\bm{r}}_{t}$, respectively.
III.1 Deterministic rate equation
The leading order $O(\Omega^{1/2})$ yields the time evolution equation for the average part ${\bm{c}}_{t}$, which can be regarded as a coarse-grained dynamics of Eq. (1):
@@ III.1 Deterministic rate equation
See Appendix A for a detailed derivation. This equation corresponds to the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which has the same form as the Lotka–Volterra equation. There are two regimes in this model (Fig. 1(b)). When $bp<d\gamma$, the stable fixed point is $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=(0,b/\gamma)$, which represents a state without zonal flows. In contrast, when $bp>d\gamma$, the stable fixed point is $((bp-d\gamma)/p^{2},d/p)$, which represents a coexistence state. Importantly, this equation does not exhibit limit cycles, contrary to the naive expectation of the existence of predator-prey oscillations. While cyclic oscillations emerge when $\gamma=0$, these cycles are not limit cycles but structurally unstable.
III.2 Linear Langevin equation
The coarse-grained description [Eqs. (3) and (4)] ignores intrinsic noise that is inevitably present in the original individual-level description (1). For a realistic situation where $\Omega$ is large but finite, it becomes essential to consider the intrinsic noise. To investigate the effect of the intrinsic noise on the coexistence state, we focus on the regime $bp>d\gamma$ and proceed to the subleading order $O(\Omega^{0})$ of the system size expansion by substituting $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=((bp-d\gamma)/p^{2},d/p)$. The resulting equation describes the time evolution of the fluctuating part ${\bm{r}}_{t}$ around ${\bm{c}}_{t}={\bm{c}}_{\mathrm{ss}}$ (see Appendix A for a detailed derivation):
@@ III.2 Linear Langevin equation
where $\tau:=bt$ denotes the dimensionless time, $\beta:=d/b$ denotes the dimensionless damping rate, and $\epsilon:=1-d\gamma/bp\in(0,1)$ denotes the ratio of time scales for $z$ and $w$. The terms $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ denote the zero-mean white Gaussian noise that satisfies $\langle\zeta^{Z}_{\tau}\zeta^{Z}_{\tau^{\prime}}\rangle=2D^{Z}\delta(\tau-\tau^{\prime})$, $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$, and $\langle\zeta^{W}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=2D^{W}\delta(\tau-\tau^{\prime})$, where $D^{Z}:=\epsilon c^{W}_{\mathrm{ss}}=\epsilon d/p$ and $D^{W}:=c^{W}_{\mathrm{ss}}=d/p$.
We emphasize that the statistical properties of the noise, such as the noise intensities $D^{Z}$ and $D^{W}$, are uniquely determined from the master equation (1). In other words, the noise driving the system is not artificially added but rather arises from intrinsic stochasticity. This property contrasts with the stochastic models proposed in Refs. [26, 27, 29, 30, 31, 32, 33], where noise is artificially added to the corresponding deterministic predator-prey models, and its properties can be arbitrarily determined. We also remark that the noises $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ in our model are not independent but are correlated as $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$, which is a property not considered in previously proposed models.
Note that the linear Langevin equations (5) and (6) have a form in which noise is added to the linearized Lotka–Volterra equation around the coexistence fixed point. To see this point explicitly, we rewrite Eqs. (5) and (6) in vector form with ${\bm{r}}_{\tau}=(z_{\tau},w_{\tau})$:
$\dfrac{d}{d\tau}{\bm{r}}_{\tau}=-\mathsf{A}{\bm{r}}_{\tau}+\mathsf{B}{\bm{\xi}}_{\tau},$ (7)
where
$\mathsf{A}=\begin{pmatrix}0&-\epsilon\\ \beta&1-\epsilon\end{pmatrix}$, (8)
and $\bm{\xi}$} denotes the zero-mean white Gaussian noise that satisfies $\langle\xi^{\alpha}_{\tau}\xi^{\alpha^{\prime}}_{\tau^{\prime}}\rangle=\delta^{\alpha\alpha^{\prime}}\delta(\tau-\tau^{\prime})$ ($\alpha=Z,W$), which is multiplied by the noise matrix $\mathsf{B}$} that satisfies
$\mathsf{B}\mathsf{B}^{\top}=\begin{pmatrix}2D^{Z}&-D^{Z}\\ -D^{Z}&2D^{W}\end{pmatrix}$. (9)
Then, it is easy to see that $\mathsf{A}$ corresponds to the coefficient matrix in the linearized Lotka–Volterra equation (3) and (4) at the coexistence fixed point.
The linear Langevin equations (5) and (6) are equivalent to the Fokker–Planck equation,
$\partial_{\tau}p_{\tau}({\bm{r}})=-\dfrac{\partial}{\partial z}J^{Z}_{\tau}({\bm{r}})-\dfrac{\partial}{\partial w}J^{W}_{\tau}({\bm{r}}),$ (10)
where $p_{\tau}({\bm{r}})$ denotes the probability density of finding ${\bm{r}}_{\tau}={\bm{r}}$ and $J^{Z}_{\tau}({\bm{r}})$ and $J^{W}_{\tau}({\bm{r}})$ denote the probability currents defined by
$J^{Z}_{\tau}({\bm{r}})$ $:=\epsilon wp_{\tau}({\bm{r}})-D^{Z}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial w}p_{\tau}({\bm{r}}),$ (11)
$J^{W}_{\tau}({\bm{r}})$ $:=\bigl[-\beta z-(1-\epsilon)w\bigr]p_{\tau}({\bm{r}})$
$\qquad-D^{W}\dfrac{\partial}{\partial w}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}}).$ (12)
Because the coefficient matrix $\mathsf{A}$ has eigenvalues $\lambda_{\pm}$ with strictly positive real parts, there exists a unique steady-state probability density $p_{\mathrm{ss}}({\bm{r}})$ for this Fokker–Planck equation [49, 50]. It can be explicitly calculated as
$p_{\mathrm{ss}}({\bm{r}})=\dfrac{1}{2\pi\sqrt{\det\Sigma}}\exp\left(-\dfrac{1}{2}({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\Sigma^{-1}({\bm{r}}-\langle{\bm{r}}\rangle)\right),$ (13)
where $\Sigma:=\langle({\bm{r}}-\langle{\bm{r}}\rangle)({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\rangle$ denotes the covariance matrix and $\langle\cdot\rangle$ denotes the average with respect to this distribution. Importantly, the steady state is generally out of equilibrium. That is, the steady-state probability current does not vanish, and there exists $\bm{r}$} such that ${\bm{J}}_{\mathrm{ss}}({\bm{r}}):=(J^{Z}_{\mathrm{ss}}({\bm{r}}),J^{W}_{\mathrm{ss}}({\bm{r}}))\neq{\bm{0}}$. In other words, the detailed balance condition (potential condition [49]) is generally violated in this linear Langevin equation. We can check that the detailed balance condition is satisfied when $\epsilon=0$.
IV Quasi-cycles induced by stochastic amplification
@@
Although the coarse-grained deterministic description [Eqs.(3) and (4)] cannot predict stable cycles, the stochastic description [Eqs. (5) and (6)] predicts persistent cyclic oscillations called quasi-cycles. Figure 1(c) shows an example of such persistent fluctuating oscillations observed in the linear Langevin equation. Notably, the fluctuations of zonal flow follow those of turbulence with a phase lag $\sim\pi/2$, which are similar to the predator-prey-like oscillations observed in a toroidal plasma [14, 16, 12, 51, 52]. Here, it should be noted that several experimental studies have also observed the opposite behavior, where turbulence follows zonal flow [15, 17, 53]. Whether such behavior can be reproduced in our model requires further investigation in future studies.
This resonant behavior can be quantified by the power spectral density matrix, which is defined as the Fourier transform of the steady-state autocorrelation function [49]:
$\mathsf{S}(\omega)$ $:=\dfrac{1}{2\pi}\int^{\infty}_{-\infty}\langle({\bm{r}}_{\tau}-\langle{\bm{r}}_{\tau}\rangle)({\bm{r}}_{0}-\langle{\bm{r}}_{0}\rangle)^{\top}\rangle e^{i\omega\tau}d\tau$
$=\dfrac{1}{2\pi}\Bigl(\mathsf{A}-i\omega\mathsf{I}\Bigr)^{-1}\mathsf{B}\mathsf{B}^{\top}\Bigl(\mathsf{A}^{\top}+i\omega\mathsf{I}\Bigr)^{-1}.$ (14)
Then, the power spectral density for zonal flow $Z$ can be calculated as
$\mathsf{S}^{ZZ}(\omega)$ $=\dfrac{C_{1}+C_{2}\omega^{2}}{(\omega^{2}-\epsilon\beta)^{2}+(1-\epsilon)^{2}\omega^{2}},$ (15)
where
$C_{1}$ $:=\dfrac{1}{\pi}\Bigl((1-3\epsilon)D^{Z}+\epsilon^{2}D^{W}\Bigr),$ (16)
$C_{2}$ $:=\dfrac{D^{Z}}{\pi}.$ (17)
From this expression, we can see that the quasi-cycle occurs for a wide range of parameters where the characteristic resonant frequency $\omega_{c}:=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}$ takes a real value (see also Fig. 2(b)). When this resonance condition is satisfied, the coexistence fixed point is a stable spiral fixed point with the characteristic frequency $|\mathrm{Im}[\lambda_{\pm}]|=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/4}$, where $\mathrm{Im}[\cdot]$ denotes the imaginary part. Note that no tuning is necessary to achieve resonance in this model. Indeed, because the system is driven by the white noise, which covers all frequencies, the resonant frequency $\omega_{c}\sim|\mathrm{Im}[\lambda_{\pm}]|$ is excited without tuning. This mechanism is called stochastic amplification to distinguish it from stochastic resonance [22]. We remark that a similar analysis can be carried out for turbulence $W$.
@@
Note that stochastic oscillations can also occur for a dynamical system with limit cycles perturbed by extrinsic noise. While quasi-cycles resemble such stochastic limit cycles, they possess several distinguishing characteristics. First, the power spectral density of quasi-cycles exhibits a long tail with $\omega^{-2}$ scaling [Eq. 15]; in contrast, stochastic limit cycles decay more rapidly as $\omega^{-4}$ [44]. This implies that the correlation time scale of fluctuations in quasi-cycles is shorter than that of stochastic limit cycles [43]. Second, the joint probability density of the oscillating variables for quasi-cycles is Gaussian [Eq. (13)], whereas that for stochastic limit cycles takes the shape of a crater ridge [43]. Furthermore, because quasi-cycles can readily emerge in any system possessing a stable spiral fixed point subject to noise, they may be more widely observed in plasma turbulence under various conditions. Although we have emphasized that our model is a qualitative description, here we present a set of parameter values consistent with experiments and simulations for reference. By comparing with the gyrokinetic simulations [54], they can be estimated as $\epsilon\sim 1$ and $\beta\sim 10^{-4}$, which satisfies the resonance condition, and the corresponding cycle frequency reads $\omega_{c}\sim 10^{-2}$ in dimensionless units.
V Information propagation between zonal flow and turbulence
The presence of quasi-cycles suggests that intrinsic fluctuations are essential for understanding the dynamics of zonal flow and turbulence. Therefore, we scrutinize the predictive causality underlying the intrinsic fluctuations from an information-theoretic viewpoint. In the context of plasma turbulence, (net) transfer entropy [24, 25] and information length [26, 27, 28, 29, 30, 31, 32, 33] have been used to quantify predictive causality. Although these quantities are useful in quantifying various aspects of causal influences, it should be noted that they are not antisymmetric under time reversal. In particular, the (net) transfer entropy can be nonzero even in an equilibrium state where all probability currents vanish [37]. (For other shortcomings of the transfer entropy, see, e.g., Refs. [55, 56].) For the information length, although it becomes zero in such an equilibrium state, it also vanishes in a nonequilibrium steady state where some probability currents remain nonzero. Therefore, from the perspective of consistency with the behavior of the probability currents, it is unclear whether these quantities appropriately quantify predictive causality. In this study, we employ information flow [34, 35, 36, 37] because it is antisymmetric under time reversal and becomes zero in the equilibrium state.